Suppose that you have a 520-Ω, a 810-Ω, and a 1.50-kΩ resistor.

What is the maximum resistance you can obtain by combining these?

What is the minimum resistance you can obtain by combining these?
Answer #1

here,

R1 = 520 ohm

R2 = 810 ohm

R3 = 1.5 Kohm = 1500 ohm

for maximum resistance , they must be connected in series

the maximum equivalent resistance , Req = R1 + R2 + R3

Req = 520 + 810 + 1500 ohm

Req = 2830 ohm

for minimum resistance , they must be connected in parallel

the minimum equivalent resistance be Req,

1/Req = 1/R1 + 1/R2 + 1/R3

1/Req = 1/520 + 1/810 + 1/1500

solving for Req

Req = 261.48 ohm